X is a normally distributed variable with mean p = 30 and standard deviation o = 4. Find P(x < 40)
Answer #1

z score for 40 = (40 - 30)/4 = 10/4 = 2.5

Hence,

P(X < 40)

= P(z < 2.5)

= 0.9938